amandak1980:

A bus is traveling with a uniform acceleration of 2.75 meters/second2. If the initial velocity of the bus is 16.5 meters/second, what will its velocity be after 9.00 seconds? A. 27.8 meters/second B. 41.3 meters/second C. 58.0 meters/second D. 408 meters/second E. 57.4 meters/second

Xaophu:

Use your kinematics equations vf=vi−at a=2.75m/s2 vi=16.5m/s t=9.00s vf=vf Plug in and solve for vf

amandak1980:

so the answer is ?

Xaophu:

I told you how to solve it

amandak1980:

I dont have nothing to solve it sorry. thats why im asking

Xaophu:

ok hold on

amandak1980:

thanks

Xaophu:

You could make a table that looks like this: Time Velocity 0 16.50 m/s (This is your starting point) 1s 19.25 m/s 2s 22.00 m/s 3s 24.75 m/s 4s 27.50 m/s 5s 30.25 m/s 6s 33.00 m/s 7s 35.75 m/s 8s 38.50 m/s 9s 41.25 m.s

amandak1980:

this are the answers: 27.8 meters/second B. 41.3 meters/second C. 58.0 meters/second D. 408 meters/second E. 57.4 meters/second

amandak1980:

that was wrong

amandak1980:

is one of them a.b.c.or d

Xaophu:

I just didnt round

Xaophu:

I mean rounded up

Xaophu:

because its 41.25 you use the 5 in .25 to make it .3

amandak1980:

i put that one. me and my son is doing plto with the books for school and i cant answer half of the questions

amandak1980:

A stone is dropped from a tower 100 meters above the ground. The stone falls past ground level and into a well. It hits the water at the bottom of the well 5.00 seconds after being dropped from the tower. Calculate the depth of the well. Given: g = -9.81 meters/second2. A. 22.5 meters B. 50.7 meters C. 100 meters D. 110 meters E. 152.45 meters

amandak1980:

A stone is dropped from a tower 100 meters above the ground. The stone falls past ground level and into a well. It hits the water at the bottom of the well 5.00 seconds after being dropped from the tower. Calculate the depth of the well. Given: g = -9.81 meters/second2. A. 22.5 meters B. 50.7 meters C. 100 meters D. 110 meters E. 152.45 meters